# KramaBench: A Benchmark for AI Systems on Data Intensive Tasks

### v2

@@ Appendix E Evaluation Details
E.1 Main metric: End-to-end automation setting
@@
| Type | Example | Metric | Scoring | ||
The name of a file to load. | Accuracy | |||
The month when an experiment started. | ParaPluie paraphrase detection (Lemesle et al., 2025) | |||
Counting the number of entries satisfying a predicate. | Accuracy | |||
Prediction of a future experiment observation. | Relative Absolute Error (RAE) | |||
Names of columns to filter data. | F1 (exact match) | F1 score | ||
Regression coefficients for different variables. | F1 score (approximate match > 0.9) | F1 score |
E.2 Additional Evaluation Settings
@@
Pipeline Design: This setting evaluates how many essential functions a system-generated pipeline includes. Here, we ask the system to provide an end-to-end pipeline implemented in Python that solves an end-to-end task. For evaluation, we manually curated an explicit list of key functionalities that any correct solution must implement for each task. We evaluate whether the generated pipeline code covers each functionality using the LLM evaluation method proposed in Tong & Zhang (2024). The score for a single task is computed as
@@
Appendix F Summary of LLM Usage
@@
LLM-as-a-judge were used to evaluade string paraphrases and code coverage. This is described in detail in Appendix E.

### v3

@@ Appendix G Evaluation Details
G.1 Main metric: End-to-end automation setting
@@
| Type | Example | Metric | Scoring |
| String (exact) | The name of a file to load. | Accuracy | |
| String (approximate) | The month when an experiment started. | ParaPluie paraphrase detection (Lemesle et al., 2025) | |
| Numeric (exact) | Counting the number of entries satisfying a predicate. | Accuracy | |
| Numeric (approximate) | Prediction of a future experiment observation. | Relative Absolute Error (RAE) | |
| List (exact) | Names of columns to filter data. | F1 (exact match) | F1 score |
| List (approximate) | Regression coefficients for different variables. | F1 score (approximate match > 0.9) | F1 score |
G.2 LLM-as-a-judge Validation
To assess the validity of the evaluation for String (approximate) and List (approximate) with String (approximate) list members conducted via instruction tuning an LLM, we performed a small scale human-LLM evaluator agreement study. We asked three human reviewers to manually evaluate the equivalence between the reference solutions and the answers generated by 12 different SUTs (the three variants of DS-Guru across four different LLM backends). We run the LLM-as-a-judge evaluation pipeline three times. We report the Cohen’s Kappa values for inter-human agreement, human-LLM agreement and inter-LLM calibration (Table 18). The possible values range from -1 (complete misalignment) to 1 (complete alignment). The results show very high inter-human agreement ( on average) and moderately high human-LLM agreement ( on average), indicating that our usage of LLM-as-a-judge provides meaningful evaluation results.
| Inter-Human Agreement | Inter-LLM Agreement | Human–LLM Agreement | ||||||
| Rater 1 | Rater 2 | Rater 1 | Rater 2 | Rater 1 | Rater 2 | |||
| Human_0 | Human_1 | 0.949 | LLM judge_0 | LLM judge_1 | 1 | Human_0 | LLM judge_0 | 0.870 |
| Human_0 | Human_2 | 0.950 | LLM judge_0 | LLM judge_2 | 1 | Human_1 | LLM judge_0 | 0.867 |
| Human_1 | Human_2 | 0.949 | LLM judge_1 | LLM judge_2 | 1 | Human_2 | LLM judge_0 | 0.818 |
G.3 Additional Evaluation Settings
@@
Pipeline Design: This setting evaluates how many essential functions a system-generated pipeline includes. Here, we ask the system to provide an end-to-end pipeline implemented in Python that solves an end-to-end task. For evaluation, we manually curated an explicit list of key functionalities that any correct solution must implement for each task. We evaluate whether the generated pipeline code covers each functionality using the LLM evaluation method proposed in Tong and Zhang (2024). The score for a single task is computed as
@@
Appendix H Summary of LLM Usage
@@
LLM-as-a-judge were used to evaluade string paraphrases and code coverage. This is described in detail in Appendix G.